# Reasoning is All You Need for Video Generalization: A Counterfactual Benchmark with Sub-question Evaluation

### v1

@@ 6 Conclusion
This paper introduces COVER, a comprehensive benchmark for counterfactual video reasoning, evaluating MLLMs across four dimensions: abstract-concrete and perception-cognition. By decomposing complex queries into structured sub-questions, COVER enables a detailed analysis of reasoning performance, highlighting the link between sub-question accuracy and overall counterfactual reasoning. Our findings emphasize that enhanced reasoning capabilities are essential for improving model performance in dynamic, real-world video tasks. COVER sets a new standard for evaluating logical reasoning in multimodal video tasks, driving the development of models that can effectively reason across dynamic video scenarios.
@@ A.1 Data Construction Details
In this section, we present additional details on COVER construction, including information about the annotation agreements, data augmentation prompts, and process flow.
@@
Three annotators were recruited to evaluate COVER across multiple dimensions, with the resultant assessments systematically compiled in Table 6, ensuring comprehensive evaluation coverage. The methodological workflow for data augmentation is schematically outlined in Figure 7.
The schematic framework outlined in Figure 5 delineates the methodology employed for contextual data augmentation, leveraging the generative capabilities of GPT-4oHurst et al. (2024) to construct domain-specific instructional prompts.

### v2

@@ 6 Conclusion
We introduce COVER, a comprehensive benchmark for counterfactual video reasoning that evaluates MLLMs across four dimensions: abstract-concrete and perception-cognition. By decomposing complex queries into structured sub-questions, COVER enables fine-grained analysis and reveals a strong correlation between sub-question accuracy and overall reasoning performance. Our results highlight the need for improved reasoning abilities in dynamic video tasks, and position COVER as a new standard for evaluating multimodal logical reasoning.
Acknowledgments
We would like to thank the anonymous reviewers for their valuable feedback. We thank Junshu Pan, Panzhong Lu, Fang Guo, Zijie Yang, Pai Liu, and other global collaborators for their valuable discussions and help. This work is funded by the National Natural Science Foundation of China Key Program (Grant No. 62336006), the Pioneer and “Leading Goose” R&D Program of Zhejiang (Grant No. 2022SDXHDX0003), and the Research Program (Grant No. WU2023C020) of the Research Center for Industries of the Future, Westlake University.
@@ A.1 Data Construction Details
In this section, we present additional details on COVER construction, including information about the task splitting scores, annotation agreements, data augmentation prompts and process flow.
We invited three expert annotators to independently score each benchmark task based on our two-dimensional quadrant framework (abstract vs. concrete and perception vs. cognition). Their scoring results in Table 8 demonstrates the strictness, consistency, and logical coherence of our task categorization, effectively preventing overlaps and ambiguity.
Task | Avgx | Avgy | ||||||
Counting | -3.2 | -3.4 | -3.1 | -3.6 | -3.3 | -3.7 | -3.2 | -3.57 |
Color | -4.1 | -4.4 | -4.4 | -4.2 | -4.2 | -4.3 | -4.23 | -4.3 |
Material | -3.8 | -3.3 | -3.9 | -3.2 | -4.0 | -3.4 | -3.9 | -3.3 |
Size | -2.4 | -2.5 | -2.6 | -2.3 | -2.2 | -2.4 | -2.4 | -2.4 |
Shape | -3.3 | -3.2 | -3.5 | -3.2 | -3.8 | -4.0 | -3.53 | -3.47 |
Emotion | -2.4 | 4.0 | -2.5 | 3.5 | -2.4 | 3.1 | -2.43 | 3.53 |
Location | -1.7 | -1.4 | -2.0 | -1.6 | -1.3 | -1.7 | -1.67 | -1.57 |
Direction | -2.1 | -1.7 | -2.5 | -1.5 | -2.6 | -1.8 | -2.4 | -1.67 |
Object Recognition | 3.0 | -3.0 | 2.4 | -2.0 | 1.2 | -2.3 | 2.2 | -2.43 |
Action Recognition | 2.5 | -3.1 | 2.3 | -3.0 | 2.1 | -3.5 | 2.3 | -3.2 |
Action Prediction | 3.9 | 2.4 | 3.8 | 2.5 | 3.2 | 2.2 | 3.63 | 2.37 |
Procedure Understanding | 3.0 | 3.5 | 3.6 | 3.2 | 2.2 | 3.3 | 2.93 | 3.33 |
Social Relation | 3.4 | 4.3 | 3.0 | 4.4 | 3.1 | 4.1 | 3.17 | 4.27 |
@@
The annotators were recruited to evaluate COVER across multiple dimensions, with the resultant assessments systematically compiled in Table 9, ensuring comprehensive evaluation coverage. The methodological workflow for data augmentation is schematically outlined in Figure 6.
@@ A.2 Additional Results
We conducted an additional ablation study to examine whether the observed trend where excessive visual information impairs reasoning accuracy holds consistently across both short and long videos. Our results are summarized in Table 10, 11. We observed a clear pattern across both short and long videos: model accuracy typically peaks within a moderate frame range (8–32 frames) and subsequently declines at the maximum setting (64 frames). This decline is particularly pronounced in tasks involving the original questions (ori) and sub questions (sub), suggesting that an excessive amount of visual input can indeed negatively impact model performance, regardless of video length.
Frame | InternVL2.5-4B | InternVL2.5-8B | ||||
2 | 69.09 | 45.94 | 60.33 | 68.72 | 56.53 | 61.14 |
4 | 70.81 | 46.18 | 60.91 | 68.97 | 56.90 | 60.68 |
8 | 71.31 | 43.97 | 60.62 | 69.83 | 56.28 | 61.14 |
16 | 70.81 | 44.83 | 59.86 | 70.07 | 56.40 | 61.20 |
32 | 70.69 | 43.84 | 59.57 | 69.21 | 56.90 | 61.26 |
64 | 69.95 | 46.55 | 59.63 | 69.09 | 57.27 | 60.62 |
Frame | InternVL2.5-4B | InternVL2.5-8B | ||||
2 | 73.90 | 49.22 | 61.09 | 72.24 | 59.26 | 59.67 |
4 | 75.79 | 46.61 | 62.13 | 75.04 | 59.07 | 61.07 |
8 | 76.46 | 46.23 | 62.24 | 75.79 | 57.37 | 61.78 |
16 | 77.45 | 45.38 | 62.31 | 75.94 | 58.27 | 61.82 |
32 | 77.40 | 45.57 | 61.49 | 75.98 | 57.08 | 61.49 |
64 | 76.13 | 46.57 | 61.11 | 76.17 | 58.41 | 61.55 |
Additionally, we evaluated test-time reasoning strategies on manually curated seed data using long-chain reasoning models in Table 12. Notably, models such as InternVL2.5-78B-CoT show significant improvement in bridging the cf–sub–ori gap, further supporting that reasoning-guided prompting (e.g., CoT) helps align sub-level and cf-level accuracy. These observations suggest a promising direction: larger and better-aligned models, when combined with explicit reasoning strategies, are more capable of maintaining coherence across perception, decomposition, and abstract reasoning tasks.
Model | ori_acc | cf_acc | sub_acc |
QVQ-72B-Preview | 69.33 | 59.33 | 58.76 |
InternVL2.5-78B-CoT | 70.00 | 71.33 | 70.80 |